Express 3 sin 4t + 8 cos 4t in the form R sin ( ωt + α ), α    


1
Expert's answer
2021-03-07T17:15:51-0500

Express 3 "\sin" 4t + 8 "\cos" 4t in the form R "\sin" ("\omega t" + "\alpha"), "\alpha"

Solution

Let 3 "\sin" 4t + 8 "\cos" 4t = R "\sin" ("\omega t" + "\alpha")

By using compound angle formula "\sin (A + B) = \sin A \cos B + \cos A \sin B" we get to solve the following:

3 "\sin" 4t + 8 "\cos" 4t = R[ "\sin" 4t "\cos \alpha" + "\cos 4t" "\sin \alpha"]

= (R "\cos \alpha") "\sin" 4t + (R "\sin \alpha") "\cos 4t"

Equating the coefficients of:

"\sin" 4t gives: 3 = R "\cos \alpha", from which, "\cos \alpha" = "\frac{3}{R}"

And "\cos" 4t gives: 8 = R "\sin \alpha", from which, "\sin \alpha" = "\frac{8}{R}"

There is only one quadrant where both "\sin \alpha" and "\cos \alpha" are positive and is the first.

Hence: R = "\sqrt{3^2 + 8^2}" = 8.544003745

from trigonometric ratios: "\alpha = \arctan \frac{8}{3}" = 69.44395478 or 1.212025657 radians

Hence 3 "\sin" 4t + 8 "\cos" 4t = 8.544003745 "\sin (4t + 1.212025657)", 1.212025657

Answer: 8.544003745 "\sin"(4t + 1.212025657), 1.212025657